Find the selling price per pound of a coffee mixture made from 8 pounds of coffee that sells for $9.20 per pound and 12 pounds of coffee that sells for $5.50 per pound.
Thank you

Answer by stanbon(75887)

Find the selling price per pound of a coffee mixture made from 8 pounds of coffee that sells for $9.20 per pound and 12 pounds of coffee that sells for $5.50 per pound.
------------
Value of 8 lbs. at 9.20= $73.60
Value of 12 lbs. at 5.50= $66
Amount of mixture = (8+12)=20 lbs.
Let "x" be the price per pound of the mixture
Value of the mixture is 20 x dollars
EQUATION:
value + value = 20x
73.60+66=20x
139.6=20x
x=$6.98 (price per pound of the mixture)
Cheers,
Stan H.